# Is Noise Conditioning Necessary for Denoising Generative Models?

### v1

@@ 1 Introduction
At the core of denoising diffusion models (Sohl-Dickstein et al., 2015) lies the idea of corrupting clean data with various levels of noise and learning to reverse this process. The remarkable success of these models has been partially underpinned by the concept of “noise conditioning(Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020): a single neural network is trained to perform denoising across all noise levels, with the noise level provided as a conditioning input. The concept of noise conditioning has been predominantly incorporated in diffusion models and is widely regarded as a critical component.
In this work, we examine the necessity of noise conditioning in denoising-based generative models. Our intuition is that, in natural data such as images, the noise level can be reliably estimated from corrupted data, making blind denoising (i.e., without knowing the noise level) a feasible task. Notably, noise-level estimation and blind image denoising have been active research topics for decades (Stahl et al., 2000; Salmeri et al., 2001; Rabie, 2005), with neural networks offering effective solutions (Chen et al., 2018; Guo et al., 2019; Zhang et al., 2023). This raises an intriguing question: can related research on image denoising be generalized to denoising-based generative models?
@@
Motivated by this, in this work, we systematically compare a variety of denoising-based generative models — with and without noise conditioning. Contrary to common belief, we find that many denoising generative models perform robustly even in the absence of noise conditioning. In this scenario, most methods exhibit only a modest degradation in generation performance. More surprisingly, we find that some relevant methods—particularly flow-based ones (Lipman et al., 2023; Liu et al., 2023), which originated from different perspectives—can even produce improved generation results without noise conditioning. Among all the popular methods we studied, only one variant fails disastrously. Overall, our empirical results reveal that noise conditioning may not be necessary for denoising generative models to function properly.
@@
Because noise-unconditional models have been rarely considered, it is worthwhile to design models specifically for this underexplored scenario. To this end, we present a simple alternative derived from the EDM model (Karras et al., 2022). Without noise conditioning, our variant can achieve a strong performance, reaching an FID of 2.23 on the CIFAR-10 dataset. This result significantly narrows the gap between a noise-unconditional system and its noise-conditional counterpart (e.g., 1.97 FID of EDM).
Looking ahead, we hope that removing noise conditioning will pave the way for new advancements in denoising-based generative modeling. For example, only in the absence of noise conditioning can a score-based model learn a unique score function and enable the classical, physics-grounded Langevin dynamics. [^1] Otherwise, it relies on the annealed Langevin dynamics (Song & Ermon, 2019)) that does not correspond to a unique underlying probability distribution independent of noise levels. . Overall, we hope that our findings will motivate the community to re-examine the fundamental principles of related methods and explore new directions in the area of denoising generative models.
@@ 2 Related Work
The seminal work of diffusion models (Sohl-Dickstein et al., 2015) proposes iteratively perturbing clean data and learning a model to reverse this process. In this pioneering work, the authors introduced a “time dependent readout function”, which is an early form of noise conditioning.
The modern implementation of noise conditioning is popularized by the introduction of Noise Conditional Score Networks (NCSN) (Song & Ermon, 2019). NCSN is originally developed for score matching. This architecture is adopted and improved in Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020), which explicitly formulate generation as an iterative denoising problem. The practice of noise conditioning has been inherited in iDDPM (Nichol & Dhariwal, 2021), ADM (Dhariwal & Nichol, 2021), and nearly all subsequent derivatives.
DDIM (Song et al., 2021a) and EDM (Karras et al., 2022) reformulate the reverse diffusion process into an ODE solver, enabling deterministic sampling from a single initial noise. Flow Matching (FM) models (Lipman et al., 2023; Liu et al., 2023; Albergo et al., 2023) reformulate and generalize the framework by learning flow fields that map one distribution to another. In all these methods, noise conditioning (also called time conditioning) is the de facto choice.
Beyond diffusion models, Consistency Models (Song et al., 2023) have emerged as a new family of generative models for non-iterative generation. It has been found (Song & Dhariwal, 2024) that noise conditioning and its implementation details are critical for the success of consistency models, highlighting the central role of noise conditioning.
@@
In the field of image processing, blind image denoising has been studied for decades. It refers to the problem of denoising an image without any prior knowledge about the level, type, or other characteristics of the noise. Relevant studies include noise level estimation from noisy images (Stahl et al., 2000; Shin et al., 2005; Liu et al., 2013; Chen et al., 2015), as well as directly learning to perform blind denoising from data (Liu et al., 2007; Chen et al., 2018; Batson & Royer, 2019; Zhang et al., 2023). Modern neural networks, including the U-Net (Ronneberger et al., 2015) commonly used in diffusion models, have been shown highly effective for these tasks.
@@ 3.1 Denoising Generative Models
where $c(t)$ and $d(t)$ are also method-specific schedule functions. Common choices of $r$ include ${\bm{\epsilon}}$-prediction (Ho et al., 2020), ${\mathbf{x}}$-prediction (Salimans & Ho, 2022), or ${\mathbf{v}}$-prediction (Salimans & Ho, 2022; Lipman et al., 2023).
The specifics of the schedule functions of several existing methods are in Table 1. It is worth noting that, in our reformulation, we concern the regression target $r$ with respect to the neural network ${\texttt{NN}}_{{\bm{\theta}}}$’s direct output.222 For methods like EDM where the network output is wrapped with a precondition, we rewrite the schedules to expose the term of ${\texttt{NN}}_{{\bm{\theta}}}$ (see Section D.3). This network ${\texttt{NN}}_{{\bm{\theta}}}$ is called the “raw network” in EDM (see Eq. (8) in (Karras et al., 2022)).
@@
     $a(t)$      $\sqrt{\bar{\alpha}(t)}$      $\frac{1}{\sqrt{t^{2}+\sigma_{\text{d}}^{2}}}$      $1-t$
@@
     $c(t)$      00      $\frac{t}{\sigma_{\text{d}}\sqrt{t^{2}+\sigma_{\text{d}}^{2}}}$      $-1$
     $d(t)$      $1$      $-\frac{\sigma_{\text{d}}}{\sqrt{t^{2}+\sigma_{\text{d}}^{2}}}$      $1$
@@ 3.2 Noise Conditional Networks
In existing methods, the neural network ${\texttt{NN}}_{{\bm{\theta}}}({\mathbf{z}}|t)$ is conditioned on the noise level specified by $t$. See Fig. 1 (left). This is commonly implemented as $t$-embedding, such as Fourier (Tancik et al., 2020) or positional embedding (Vaswani, 2017). This $t$-embedding provides time-level information as an additional input to the network. Our study concerns the influence of this noise conditioning, that is, we consider ${\texttt{NN}}_{{\bm{\theta}}}({\mathbf{z}})$ vs. ${\texttt{NN}}_{{\bm{\theta}}}({\mathbf{z}}|t)$. See Fig. 1 (right). Note that ${\texttt{NN}}_{{\bm{\theta}}}({\mathbf{z}})$ or ${\texttt{NN}}_{{\bm{\theta}}}({\mathbf{z}}|t)$ involves all learnable parameters in the model, while the schedules ($a(t)$, $b(t)$, etc.) are pre-designed and not learned.
@@ 4 Analysis of Noise-Unconditional Models
Figure 2: A given $z$ corresponds to multiple triplets $({\mathbf{x}},{\bm{\epsilon}},t)$. Here, we take Flow Matching (Lipman et al., 2023) as an example. On the left are the samples of ${\bm{\epsilon}}$, and on the right are samples of ${\mathbf{x}}$. For a noisy sample ${\mathbf{z}}=(1-t){\mathbf{x}}+t{\bm{\epsilon}}$, it can be produced by different triplets. Each triplet gives a different regression target $r$. The effective target $R({\mathbf{z}})$ is the expectation of all possible $r$.
@@ 4.1 Effective Targets
Here, $p({\mathbf{z}})$ is the marginalized distribution of ${\mathbf{z}}{:=}a(t){\mathbf{x}}+b(t){\bm{\epsilon}}$ in Eq. 1, under the joint distribution $p({\mathbf{x}},{\bm{\epsilon}},t):=p({\mathbf{x}})p({\bm{\epsilon}})p(t)$. [^3] For simplicity, we consider $w(t){=}1$, which happens to be the case for all methods in Table 1 when we expose ${\texttt{NN}}_{{\bm{\theta}}}$ explicitly. It is easy to show that:
@@
that is, the expectation over all $({\mathbf{x}},{\bm{\epsilon}})$ subject to the conditional distribution. One can show (Section C.1) that minimizing Eq. 5 is equivalent to minimizing Eq. 2.
@@
Next, we investigate how similar $p(t|{\mathbf{z}})$ is to a Dirac delta function. For high-dimensional data such as images, it has been long realized that the noise level can be reliably estimated (Stahl et al., 2000; Salmeri et al., 2001; Shin et al., 2005), implying a concentrated $p(t|{\mathbf{z}})$. We note that the concentration of $p(t|{\mathbf{z}})$ depends on data dimensionality:
Figure 3: Posterior distribution $p(t|{\mathbf{z}})$ given different noisy images ${\mathbf{z}}=(1-t_{*}){\mathbf{x}}+t_{*}{\bm{\epsilon}}$, for $t_{*}$ from 0.1 to 0.9. This plot is empirically simulated from 15,000 images in the AFHQ-v2 dataset with a size $64{\times}64$ (see Section A.2).
@@
Intuitively, this statement suggests that high-dimensional data will lead to $p(t|{\mathbf{z}})$ of low variance. Because this analysis is based on simplified assumptions, we empirically run a simulation on a real dataset and plot $p(t|{\mathbf{z}})$ (see Fig. 3). The empirical distribution of $p(t|{\mathbf{z}})$ is well concentrated. Moreover, a smaller $t^{*}$ leads to a more concentrated $p(t|{\mathbf{z}})$, as also indicated by Eq. 9.
@@ 4.4 Accumulated Error in Sampling
$\vspace{-1em}A_{i}=\prod_{j=i+1}^{N-1}(\kappa_{i}+|\eta_{i}|L_{i})\quad\text{and}\quad B_{i}=|\eta_{i}|\delta_{i}.$
@@ 5 A Noise Unconditional Diffusion Model
In addition to investigating existing models, we also design a diffusion model specifically tailored for noise unconditioning. Our motivation is to find schedule functions that are more robust in the absence of noise conditioning, while still maintaining competitive performance. To this end, we build upon the highly effective EDM framework (Karras et al., 2022) and modify its schedules.
@@
$c_{\text{skip}}(t)\hat{{\mathbf{z}}}+c_{\text{out}}(t){\texttt{NN}}_{{\bm{\theta}}}\big{(}c_{\text{in}}(t)\hat{{\mathbf{z}}}\mid t\big{)}\vspace{-1em}$
Here, $\hat{{\mathbf{z}}}:={\mathbf{x}}+t{\bm{\epsilon}}$ is the noisy input before the normalization performed by $c_{\text{in}}(t)$, [^4] To make notations consistent with our reformulation in Eq. 2, we denote ${\mathbf{z}}=c_{\text{in}}(t)\hat{{\mathbf{z}}}$. See details in Section D.3. which we simply set as $c_{\text{in}}(t)=\frac{1}{\sqrt{1+t^{2}}}$. The main modification we adopt for the noise unconditioning scenario is to set:
@@ 6 Experiments
  • Diffusion: iDDPM (Nichol & Dhariwal, 2021), DDIM (Song et al., 2021a), ADM (Dhariwal & Nichol, 2021), EDM (Karras et al., 2022), and uEDM (Sec. 5)
  • Flow-based Models: we adopt the implementation of Rectified Flow (1-RF) (Liu et al., 2023), which is a form of Flow Matching (Lipman et al., 2023) (FM).
  • Consistency Models: iCT (Song & Dhariwal, 2024) and ECM (Geng et al., 2025).
Our main experiments are on class-unconditional generation on CIFAR-10 (Krizhevsky et al., 2009), with extra results on ImageNet $32{\times}32$ (Deng et al., 2009), and FFHQ $64{\times}64$ (Karras et al., 2019). We evaluate Fréchet Inception Distance (FID) (Heusel et al., 2017) and report Number of Function Evaluations (NFE). For a fair comparison, all methods are based on our re-implementation as faithful as possible (see Section B.3): with and without noise conditioning are run in the same implementation for each method.
@@ 6.1 Main Observations
Table 2 summarizes the FID changes in different generative models, with and without noise conditioning, denoted as “w/ $t$” and “w/o $t$”. Fig. 5 shows some qualitative results. We draw the following observations:
@@
(ii) More surprisingly, some flow-based variants can achieve improved FID (green) after removing noise conditioning. In general, flow-based methods investigated in this paper are insensitive to whether we use noise conditioning or not. We hypothesize that this is partially because FM’s regression target is independent on $t$ (see Table 1: $c=-1$, $d=1$)
(iii) The uEDM variant (Sec. 5) achieves a competitive FID of 2.23 without noise conditioning, narrowing the gap to the strong baseline of the noise-conditional methods (here, 1.99 of EDM, or 1.97 reported in Karras et al. (2022)).
(iv) Consistency Models (here, iCT and ECM), which are related to diffusion models but present a substantially different objective function, can also perform gracefully. While iCT was found highly sensitive to the subtleties of $t$-conditioning (see Song & Dhariwal (2024)), we find that removing it does not lead to disastrous failure.
@@
Model accum. bound FID w/ $t$rightarrow w/o $t$
@@
Figure 5: Samples of noise-conditional vs. noise-unconditional models. Samples are generated by (a) DDIM, (b) EDM, (c) FM (1-RF), and (d) uEDM, on the CIFAR-10 class-unconditional case. For each subfigure, the left panel is the noise-conditional case, and the right panel is the noise-unconditional counterpart, with the same random seeds. The change of FID is from “w/ $t$” to “w/o $t$”. See also Table 2 for more quantitative results.

### v2

@@ 1 Introduction
At the core of denoising diffusion models (sohl2015diffusion) lies the idea of corrupting clean data with various levels of noise and learning to reverse this process. The remarkable success of these models has been partially underpinned by the concept of “noise conditioning(sohl2015diffusion; song2019ncsn; ho2020denoising): a single neural network is trained to perform denoising across all noise levels, with the noise level provided as a conditioning input. The concept of noise conditioning has been predominantly incorporated in diffusion models and is widely regarded as a critical component.
In this work, we examine the necessity of noise conditioning in denoising-based generative models. Our intuition is that, in natural data such as images, the noise level can be reliably estimated from corrupted data, making blind denoising (i.e., without knowing the noise level) a feasible task. Notably, noise-level estimation and blind image denoising have been active research topics for decades (stahl2000quantile; salmeri2001noise; rabie2005robust), with neural networks offering effective solutions (chen2018image; guo2019toward; zhang2023blind). This raises an intriguing question: can related research on image denoising be generalized to denoising-based generative models?
@@
Motivated by this, in this work, we systematically compare a variety of denoising-based generative models — with and without noise conditioning. Contrary to common belief, we find that many denoising generative models perform robustly even in the absence of noise conditioning. In this scenario, most methods exhibit only a modest degradation in generation performance. More surprisingly, we find that some relevant methods—particularly flow-based ones (lipman2023flow; liu2023flow), which originated from different perspectives—can even produce improved generation results without noise conditioning. Among all the popular methods we studied, only one variant fails disastrously. Overall, our empirical results reveal that noise conditioning may not be necessary for denoising generative models to function properly.
@@
Because noise-unconditional models have been rarely considered, it is worthwhile to design models specifically for this underexplored scenario. To this end, we present a simple alternative derived from the EDM model (karras2022edm). Without noise conditioning, our variant can achieve a strong performance, reaching an FID of 2.23 on the CIFAR-10 dataset. This result significantly narrows the gap between a noise-unconditional system and its noise-conditional counterpart (e.g., 1.97 FID of EDM).
Looking ahead, we hope that removing noise conditioning will pave the way for new advancements in denoising-based generative modeling. For example, only in the absence of noise conditioning can a score-based model learn a unique score function and enable the classical, physics-grounded Langevin dynamics. [^1] Otherwise, it relies on the annealed Langevin dynamics (song2019ncsn)) that does not correspond to a unique underlying probability distribution independent of noise levels. Overall, we hope that our findings will motivate the community to re-examine the fundamental principles of related methods and explore new directions in the area of denoising generative models.
@@ 2 Related Work
The seminal work of diffusion models (sohl2015diffusion) proposes iteratively perturbing clean data and learning a model to reverse this process. In this pioneering work, the authors introduced a “time dependent readout function”, which is an early form of noise conditioning.
The modern implementation of noise conditioning is popularized by the introduction of Noise Conditional Score Networks (NCSN) (song2019ncsn). NCSN is originally developed for score matching. This architecture is adopted and improved in Denoising Diffusion Probabilistic Models (DDPM) (ho2020denoising), which explicitly formulate generation as an iterative denoising problem. The practice of noise conditioning has been inherited in iDDPM (nichol2021iddpm), ADM (dhariwal2021diffusion), and nearly all subsequent derivatives.
DDIM (song2021ddim) and EDM (karras2022edm) reformulate the reverse diffusion process into an ODE solver, enabling deterministic sampling from a single initial noise. Flow Matching (FM) models (lipman2023flow; liu2023flow; albergo2023stochastic) reformulate and generalize the framework by learning flow fields that map one distribution to another. In all these methods, noise conditioning (also called time conditioning) is the de facto choice.
Beyond diffusion models, Consistency Models (song2023consistency) have emerged as a new family of generative models for non-iterative generation. It has been found (song2024improved) that noise conditioning and its implementation details are critical for the success of consistency models, highlighting the central role of noise conditioning.
@@
In the field of image processing, blind image denoising has been studied for decades. It refers to the problem of denoising an image without any prior knowledge about the level, type, or other characteristics of the noise. Relevant studies include noise level estimation from noisy images (stahl2000quantile; shin2005block; liu2013single; chen2015efficient), as well as directly learning to perform blind denoising from data (liu2007automatic; chen2018image; batson2019noise2self; zhang2023blind). Modern neural networks, including the U-Net (ronneberger2015u) commonly used in diffusion models, have been shown highly effective for these tasks.
@@ 3.1 Denoising Generative Models
where $c(t)$ and $d(t)$ are also method-specific schedule functions. Common choices of $r$ include $\bm{\epsilon}$}-prediction (ho2020denoising), $\mathbf{x}$}-prediction (salimans2022progressive), or $\mathbf{v}$}-prediction (salimans2022progressive; lipman2023flow).
The specifics of the schedule functions of several existing methods are in Table 1. It is worth noting that, in our reformulation, we concern the regression target $r$ with respect to the neural network ${\texttt{NN}}_{{\bm{\theta}}}$’s direct output. [^2] For methods like EDM where the network output is wrapped with a precondition, we rewrite the schedules to expose the term of ${\texttt{NN}}_{{\bm{\theta}}}$ (see Section D.3). This network ${\texttt{NN}}_{{\bm{\theta}}}$ is called the “raw network” in EDM (see Eq. (8) in (karras2022edm)).
@@
   $a(t)$    $\sqrt{\bar{\alpha}(t)}$    $\frac{1}{\sqrt{t^{2}+\sigma_{\text{d}}^{2}}}$    1t1-t
@@
   $c(t)$    0    $\frac{t}{\sigma_{\text{d}}\sqrt{t^{2}+\sigma_{\text{d}}^{2}}}$    1-1
   $d(t)$    11    $\frac{\sigma_{\text{d}}}{\sqrt{t^{2}+\sigma_{\text{d}}^{2}}}$    11
@@ 3.2 Noise Conditional Networks
In existing methods, the neural network ${\texttt{NN}}_{{\bm{\theta}}}({\mathbf{z}}|t)$ is conditioned on the noise level specified by $t$. See Fig. 1 (left). This is commonly implemented as $t$-embedding, such as Fourier (tancik2020fourier) or positional embedding (vaswani2017attention). This $t$-embedding provides time-level information as an additional input to the network. Our study concerns the influence of this noise conditioning, that is, we consider ${\texttt{NN}}_{{\bm{\theta}}}({\mathbf{z}})$ vs. ${\texttt{NN}}_{{\bm{\theta}}}({\mathbf{z}}|t)$. See Fig. 1 (right). Note that ${\texttt{NN}}_{{\bm{\theta}}}({\mathbf{z}})$ or ${\texttt{NN}}_{{\bm{\theta}}}({\mathbf{z}}|t)$ involves all learnable parameters in the model, while the schedules ($a(t)$, $b(t)$, etc.) are pre-designed and not learned.
@@ 4 Analysis of Noise-Unconditional Models
Figure 2: Illustration of the effective target $R({\mathbf{z}})$. A given $z$ corresponds to multiple triplets $({\mathbf{x}},{\bm{\epsilon}},t)$. Here, we take Flow Matching (lipman2023flow) as an example. On the left are the samples of $\bm{\epsilon}$}, and on the right are samples of $\mathbf{x}$}. For a noisy sample ${\mathbf{z}}=(1-t){\mathbf{x}}+t{\bm{\epsilon}}$, it can be produced by different triplets. Each triplet gives a different regression target $r$. The effective target $R({\mathbf{z}})$ is the expectation of all possible $r$.
@@ 4.1 Effective Targets
Here, $p({\mathbf{z}})$ is the marginalized distribution of ${\mathbf{z}}{:=}a(t){\mathbf{x}}+b(t){\bm{\epsilon}}$ in Eq. 1, under the joint distribution $p({\mathbf{x}},{\bm{\epsilon}},t):=p({\mathbf{x}})p({\bm{\epsilon}})p(t)$. [^3] For simplicity, we consider w(t)=1w(t){=}1, which happens to be the case for all methods in Table 1 when we expose ${\texttt{NN}}_{{\bm{\theta}}}$ explicitly. It is easy to show that:
@@
that is, the expectation over all $({\mathbf{x}},{\bm{\epsilon}})$ subject to the conditional distribution. One can show (Section C.1) that minimizing Eq. 5 is equivalent to minimizing Eq. 2, and similar analysis has also been done in previous work(lehtinen2018noise2noise).
@@
Next, we investigate how similar $p(t|{\mathbf{z}})$ is to a Dirac delta function. For high-dimensional data such as images, it has been long realized that the noise level can be reliably estimated (stahl2000quantile; salmeri2001noise; shin2005block), implying a concentrated $p(t|{\mathbf{z}})$. We note that the concentration of $p(t|{\mathbf{z}})$ depends on data dimensionality:
Figure 3: The Posterior distribution $p(t|{\mathbf{z}})$ is concentrated. We picked ${\mathbf{z}}=(1-t_{*}){\mathbf{x}}+t_{*}{\bm{\epsilon}}$ with $t_{*}$ from 0.1 to 0.9 for illustration. This plot is empirically simulated from 15,000 images in the AFHQ-v2 dataset with a size $64{\times}64$ (see Section A.2).
@@
Intuitively, this statement suggests that high-dimensional data induces a sharply peaked $p(t\mid{\mathbf{z}})$. In Section C.2, we derive a rigorous upper bound on this variance and extend the analysis to the multi–data–point setting. To corroborate these theoretical findings, we empirically run a simulation on a real dataset and plot $p(t|{\mathbf{z}})$ (see Fig. 3). The empirical distribution of $p(t|{\mathbf{z}})$ is well concentrated. Moreover, a smaller $t^{*}$ leads to a more concentrated $p(t|{\mathbf{z}})$, as also indicated by Eq. 9.
@@ 4.4 Accumulated Error in Sampling
$\vskip-10.00002ptA_{i}=\prod_{j=i+1}^{N-1}(\kappa_{i}+|\eta_{i}|L_{i})\quad\text{and}\quad B_{i}=|\eta_{i}|\delta_{i}.$
@@ 5 A Noise Unconditional Diffusion Model
In addition to investigating existing models, we also design a diffusion model specifically tailored for noise unconditioning. Our motivation is to find schedule functions that are more robust in the absence of noise conditioning, while still maintaining competitive performance. To this end, we build upon the highly effective EDM framework (karras2022edm) and modify its schedules.
@@
$c_{\text{skip}}(t)\hat{{\mathbf{z}}}+c_{\text{out}}(t){\texttt{NN}}_{{\bm{\theta}}}\big(c_{\text{in}}(t)\hat{{\mathbf{z}}}\mid t\big)\vskip-10.00002pt$
Here, $\hat{{\mathbf{z}}}:={\mathbf{x}}+t{\bm{\epsilon}}$ is the noisy input before the normalization performed by $c_{\text{in}}(t)$, [^4] To make notations consistent with our reformulation in Eq. 2, we denote $^{\mathbf{z}}=c_{\text{in}}(t)\hat{{\mathbf{z}}}$. See details in Section D.3. which we simply set as $c_{\text{in}}(t)=\frac{1}{\sqrt{1+t^{2}}}$. The main modification we adopt for the noise unconditioning scenario is to set:
@@ 6 Experiments
  • Diffusion: iDDPM (nichol2021iddpm), DDIM (song2021ddim), ADM (dhariwal2021diffusion), EDM (karras2022edm), and uEDM (Sec. 5)
  • Flow-based Models: we adopt the implementation of Rectified Flow (1-RF) (liu2023flow), which is a form of Flow Matching (lipman2023flow) (FM).
  • Consistency Models: iCT (song2024improved) and ECM (geng2025consistency).
Our main experiments are on class-unconditional generation on CIFAR-10 (krizhevsky2009CIFAR), with extra results on ImageNet ${\times}32$ (deng2009imagenet), and FFHQ ${\times}64$ (karras2019style). We evaluate Fréchet Inception Distance (FID) (heusel2017FID) and report Number of Function Evaluations (NFE). For a fair comparison, all methods are based on our re-implementation as faithful as possible (see Section B.3): with and without noise conditioning are run in the same implementation for each method.
@@ 6.1 Main Observations
Table 2 summarizes the FID changes in different generative models, with and without noise conditioning, denoted as “w/ tt” and “w/o tt”. Fig. 5 shows some qualitative results. We draw the following observations:
@@
(ii) More surprisingly, some flow-based variants can achieve improved FID (green) after removing noise conditioning. In general, flow-based methods investigated in this paper are insensitive to whether we use noise conditioning or not. We hypothesize that this is partially because FM’s regression target is independent on $t$ (see Table 1: c=1c=-1, $d=1$)
(iii) The uEDM variant (Sec. 5) achieves a competitive FID of 2.23 without noise conditioning, narrowing the gap to the strong baseline of the noise-conditional methods (here, 1.99 of EDM, or 1.97 reported in karras2022edm).
(iv) Consistency Models (here, iCT and ECM), which are related to diffusion models but present a substantially different objective function, can also perform gracefully. While iCT was found highly sensitive to the subtleties of $t$-conditioning (see song2024improved), we find that removing it does not lead to disastrous failure.
@@
Model accum. bound FID w/ $t$ $\rightarrow$ w/o $t$
@@
Figure 5: Samples of noise-conditional vs. noise-unconditional models. Samples are generated by (a) DDIM, (b) EDM, (c) FM (1-RF), and (d) uEDM, on the CIFAR-10 class-unconditional case. For each subfigure, the left panel is the noise-conditional case, and the right panel is the noise-unconditional counterpart, with the same random seeds. The change of FID is from “w/ tt” to “w/o tt”. See also Table 2 for more quantitative results.